Q 15: An electric spark jumps from a person's finger to your nose. While passing through...

Q 15:

An electric spark jumps from a person's finger to your nose. While passing through the air, the spark travels across a potential difference of 2.0×104V and releases 2.8×10−7J of electric potential energy.

Part A:

What is the charge in coulombs?

Express your answer in coulombs.

Part B:

How many electrons flow?

Express your answer numerically.

Q 16:

Two −2.4×10−9−2.4×10−9-C charged point-like objects are separated by 0.20 m.

Part A:

Determine the potential (assuming zero volts at infinity) at a point halfway between the objects.

Express your answer with the appropriate units.

Part B:

Determine the potential (assuming zero volts at infinity) at a point 0.20 m to the side of one of the objects (and 0.40 m from the other) along a line joining them.

Express your answer with the appropriate units.
Answer #1

Q15:

Relation between electric potential energy and potential difference is given by:

dU = q*dV

dU = electric potential energy = 2.8*10^-7 J

dV = potential difference = 2.0*10^4 V

So,

q = dU/dV

q = 2.8*10^-7/(2.0*10^4)

q = 1.4*10^-11 C = Charge flow

Part B.

Number of electrons flow will be given by:

q = |n|*e

|n| = number of electrons = q/e

e = charge on single electron = -1.6*10^-19

n = q/e

n = 1.4*10^-11/(1.6*10^-19)

n = 8.75*10^7 = number of electrons flow

As per guideline please ask one question at a time.